A steam turbine receives 362,500 kg/h of steam throttled at 8.7 MPa and 410OC. steam leaves the high pressure turbine shell at 2 MPa with 7.58 degrees of superheat. Reheated steam is at 1.9 MPa and 410OC. Condensate is at 78OC. For a generator output of 100 MW at 94% efficiency, calculate (a) the combined thermal efficiency, (b) the actual condition of exhaust steam, (c) the thermal efficiency of the ideal engine, (d) the combined engine efficiency, and (e) the ideal thermal efficiency of the cycle without pressure drop through reheater


1
Expert's answer
2021-10-28T07:18:57-0400

mass rate = 362,500 kg/hr

Pe = 8.7 MPa

Te = 410°C = 410 + 273 = 683K

Po = 2MPa

To = 758°C

Pr = 1.9MPa

Tr = 410°C

Thermal Efficiency = (758-410)/758 = 45,9 %